OpenStudy (anonymous):

Shelly sews a blanket that has an area of 170 square feet. It has 30 square blocks, each the same size. What is the approximate length of each side of a block?

OpenStudy (anonymous):

\[Area_{blanket} = 30Area_{\square}\] \[170 = 30Area_{\square}\]\[Area_{sqrare}=17/3\]\[b^{2}=17/3\]\[b=2.380...\]\[b {\approx}2.38\]